OpenStudy (anonymous):

how do i find the derivative of f(x) = e^(1-2x^2) + ln (3-e^-4x)?

OpenStudy (anonymous):

i need the answer...my exam is tomorrow! please help :(

OpenStudy (anonymous):

ok i can help u

OpenStudy (anonymous):

do u need the method or just theanswer

OpenStudy (anonymous):

the method please

OpenStudy (anonymous):

ok here it goes d/dx(e^(1-2x^2)) is -4xe^(1-2x^2) because d/dx(e^f(x)) is e^(x)*d/dx(f(x)

OpenStudy (anonymous):

here f(x) is 1-2x^2 therefore d/dx(1-2x^2) is -4x

OpenStudy (anonymous):

then d/dx(ln(3-e^-4x) is 4e^-4x/(3-e^-4x)

OpenStudy (anonymous):

because d/dx(log(f(x)) is (1/f(x))*d/dxf(x)

OpenStudy (anonymous):

where f(x) is (3-e^-4x)

OpenStudy (anonymous):

so the answer is -4xe^(1-2x^2)+4e^-4x/(3-e^-4x)

OpenStudy (anonymous):

what theorem or rules apply to this type of problem?

OpenStudy (anonymous):

first of all d/dx(a+b) is da/dx+db/dx where a is e^(1-2x^2) and b is ln(3-e^-4x

OpenStudy (anonymous):

first of all d/dx(a+b) is da/dx+db/dx where a is e^(1-2x^2) and b is ln(3-e^-4x)

OpenStudy (anonymous):

Maybe this will help: the derivative of e ^ (stuff) = e ^ (stuff) * (derivative of that stuff) the derivative of ln (stuff) = 1/(stuff) * (derivative of that stuff)

OpenStudy (anonymous):

yes exactly

OpenStudy (anonymous):

And as dipin said, if you have the derivative of things added or subtracted, you just take each derivative as you come across it. So derivative of (A + B) = derivative of A + derivative of B

OpenStudy (anonymous):

It's just a big ugly when you have to type these things instead of write them by hand. :)

OpenStudy (anonymous):

so we just solve the e^stuff and ln^stuff and then add them?

OpenStudy (anonymous):

yes,that's ll

OpenStudy (anonymous):

i just get confused because of the logs, e's, and ln's.

OpenStudy (anonymous):

If you can just try to step back and see the overall structure of the question (ignoring the specific bits), sometimes that helps

OpenStudy (anonymous):

one thing u have to remember is d/dx of e^x is e^x and d/dx of log x is 1/x

OpenStudy (anonymous):

Once you know the derivative of e^(stuff) = e^(stuff) * (derivative of that stuff), then you can make the stuff as complicated as you want, and you still have a basic framework for solving the question. (Of course, assuming all the proper variables like the derivative we're taking is dy/dx and the function is expressed in terms of x (or x and y, if you're there yet))

OpenStudy (anonymous):

for example e^4x is e^4x*d/dx(4x) which is 4e^4x

OpenStudy (anonymous):

You don't even have to remember d/dx e^x = e^x, if you don't want to, because if you follow the pattern, you'll get there anyway. e^x just the special case where the (stuff) is x. And the derivative of x = 1. So, you get e^x * 1 when you take the derivative with respect to x. That's the nice thing about following the rules. :)

OpenStudy (anonymous):

And then, you just clean it up like dipin does, usually writing it cleanly and compactly

OpenStudy (anonymous):

d/dx(1-2x^2) will be -2*2*x which is -4x and we have a e^(stuff) there

OpenStudy (anonymous):

oh! can i still stuff you guys into my head during the exam? :)

OpenStudy (anonymous):

anyways gud luck for ur exam.

OpenStudy (anonymous):

haha thanks for everything!

OpenStudy (anonymous):

Yes, good luck!

OpenStudy (anonymous):

thank you